OpenStudy (anonymous):

If you know plz tell me: In ideal gas how we get integral of [e^-(βp²/2m) dp] with limit of (-∞ to ∞) is = (2mπ/β)^½

OpenStudy (festinger):

I am too lazy to type the β/2m is i'll let it be k. If you want a better treatment of this integral, look for gaussian integral. \[Let\:\int_{-\infty}^{\infty}e^{-ap^{2}}dp=I\]Square both sides \[\int_{-\infty}^{\infty}\int_{\infty}^{-\infty}e^{-ap^{2}}e^{-aq^{2}}dpdq=I^{2}=\int_{\infty}^{-\infty}\int_{\infty}^{-\infty}e^{-a(p^{2}+q^{2})}dpdq\] Recall that x^2+y^2 represents a circle with radius R, so \[I^2=\int_{2\pi}^{0}\int_{0}^{\infty}e^{-ar^{2}}rdrd\theta\] \[I^2=2\pi\int_{0}^{\infty}e^{-ar^{2}}\frac{1}{2}(dr^{2})\] \[I^2=2\pi\int_{0}^{\infty}e^{-ar^{2}}\frac{1}{2}(dr^{2})\] \[I^2=0-(-\frac{\pi}{a})\] Thus,\[I=\sqrt{\frac{\pi}{a}}\]